# JWST imaging of the Pleiades: anisotropy of turbulence in the cold neutral medium

### v1

@@
Interstellar medium studies rely on magnetohydrodynamic (MHD) turbulence as a framework for interpretation. In this context, the statistical characterization of interstellar observations is of prime importance. We open a new perspective on diffuse interstellar matter by analyzing James Webb Space Telescope (JWST) observations of the Pleiades nebula with NIRCam. These observations are remarkable in that they provide a microscope view at the cold neutral medium (CNM) with a spatial resolution of 0.2 mpc (40 au). A two-dimensional Fourier analysis is used to characterize the structure of PAH emission in regions near and far from the Pleiades star Merope. To produce maps of the interstellar emission, stars and galaxies are filtered out. The final step in the data cleaning involves subtracting a component, in Fourier space, which we infer to be a residual of the near-infrared cosmic background. The PAH emission power spectra are highly anisotropic. They are well fitted with a break-free power-law, suggesting that we do not observe a specific scale for energy dissipation. Power-law indices are -3.5 near Merope and -3 in the more distant field. The magnetic field orientation, as derived from the Planck dust polarization data, aligns with the PAH anisotropy. The power anisotropy is constant across scales. These findings are discussed in relation to interstellar turbulence that may be driven by the Pleiades stars. The JWST observations of the Pleiades offer a new viewpoint for comparing observations and theoretical models, as they examine physical scales at which turbulence in the CNM is subsonic and decoupled from the thermal instability. The observations may indicate that the turbulent energy cascade in the CNM is anisotropic.
@@
Studies of the cold interstellar medium (ISM) frequently rely on numerical simulations of magnetohydrodynamic (MHD) turbulence as a theoretical model (Hennebelle and Falgarone, 2012; Burkhart, 2021). In this context, the statistical analysis of interstellar observations is crucial to interpret observations (Richard et al., 2025). This paper opens a new perspective, using observations of the Pleiades nebula by the James Webb Space Telescope (JWST) with the Near-Infrared Camera (NIRCam).
@@
The Pleiades reflection nebula is an iconic astronomical object familiar to a wide audience and also a prime scientific target to study the dusty interstellar medium (Gibson and Nordsieck, 2003a, b; Gibson, 2007). We are observing a chance encounter between a star cluster and diffuse molecular gas (Gordon and Arny, 1984; Herbig and Simon, 2001; Ritchey et al., 2006). The brightest members of the Pleiades cluster are late B-stars at a distance of 135 pc (Abramson, 2018), illuminating their surrounding ISM over a few parsecs (Gibson and Nordsieck, 2003a). The stellar illumination enables the ISM structure to be observed at smaller angular scales than is usually possible. The reflection nebulosity offers a spectacular view at the filamentary structure of the diffuse ISM.
@@
Many studies have relied on dust continuum observations to determine the power spectrum of interstellar matter (Gautier et al., 1992; Ingalls et al., 2004; Miville-Deschênes et al., 2007, 2010, 2016). The advent of Planck made it possible to extend this type of analysis to polarized dust emission (Planck Collaboration, 2020a), which established the statistical alignment of diffuse interstellar structures with the magnetic field direction on parsec scales (Clark et al., 2015; Planck Collaboration, 2016c). This result has been interpreted as a hallmark of magnetized interstellar turbulence (Caldwell et al., 2017; Kandel et al., 2017). In a broader context, the James Webb Space Telescope (JWST) offers a unique opportunity. Its angular resolution and sensitivity to the emission of polycyclic aromatic hydrocarbons (PAHs) enable imaging of the cold interstellar medium (ISM) at finer scales than was previously possible. At the distance of the Pleiades, the angular resolution of NIRCam corresponds to a spatial scale of 0.2 mpc, at which we may be able to identify signatures of dissipation of turbulence, in particular through ion-neutral friction (Burkhart et al., 2015; Hu et al., 2024). Although there is observational evidence of structure in the cold neutral medium (CNM) at these scales, existing observations are too sparse to conduct a statistical study (Heiles, 1997; Stanimirović and Zweibel, 2018).
@@
The purpose of this work is to demonstrate and capitalize on the potential of JWST for diffuse ISM studies. We use PAH emission to investigate the structure of the cold magnetized ISM. In fact, PAHs are tightly coupled to the gas through collisions (Omont, 1986) and to magnetic field lines. For physical conditions in the Pleiades nebula – specifically the gas density and temperature, and the radiation field strength – PAHs exist partly as cations and partly in neutral form (Flagey et al., 2006). However, both charged and neutral PAHs are coupled to magnetic field lines because their charge state varies on timescales much shorter than their drag time through collisions with the gas (Lepp and Dalgarno, 1988).
@@
The paper is organized as follows. Section 2 introduces the JWST observations and outlines the data processing steps to generate images of diffuse emission. Our two-dimensional Fourier analysis of these images is described in Sect. 3 and Sect. 4. In Sect. 5, we discuss the results of the data analysis in relation to interstellar turbulence. Finally, Section 6 summarizes our main results. The paper includes three appendices. Appendix A details and illustrates the methods used to remove stars and mask galaxies from the JWST images. Appendix B provides a broader context for the JWST observations, in which we characterize the magnetized interstellar matter associated with the Pleiades nebula using Planck data. In Appendix C, we use the observations to characterize the physical conditions and demonstrate that the JWST observations of the Pleiades probe the structure of the cold neutral medium.
@@ 2.1 JWST Observations
The NIRCam instrument (Rieke et al., 2023) was used to image two fields, each measuring by as part of the GO 2143 program. Field 1 is centered approximately 2′ to the East of the bright Pleiades star Merope (23 Tau), and Field 2 about 10′ to the South-East of Merope. The top images in Fig. 1 show the location of the two fields within the Pleiades. The two images at the bottom were produced by the JWST pipeline. For this paper, we used version 1.18.0 of the calibration software.
The two fields were imaged in the F212N filter, centered on the 1-0 S(1) H2 line, and the F335M filter that includes the 3.3-m emission bands from polycyclic aromatic hydrocarbons (PAHs). We also obtained observations at both sky positions in the F200W and F277W filters. The observations in the short- and long wavelength channels are executed simultaneously, in the spectral filters F212N with F335M and F200W with F277W. For each pointing, the integration time was for the H2 and PAH observations, and for the wide filters. In this study, we only analyze the F335M data. The data obtained using the three additional filters will be presented in a future paper.
The long channel of NIRCam comprises two square detectors, each with 2000 pixels square and a field of view of . We used the FULLBOX 6TIGHT111NIRCam page at https://jwst-docs.stsci.edu/jwst-near-infrared-camera/nircam-observing-modes/ dithering pattern to obtain multiple offset pointings of the camera and cover the rectangular field without gaps. The dithering also mitigates bad pixels and flat-field uncertainties; it significantly reduces the data noise on the smallest angular scales, where brightness variations are the smallest. We did not use secondary subpixel dithers because the fine sampling of the point-spread function is not essential to our observations.
@@ 2.2 Maps of diffuse emission
In order to study the PAH emission, we need to separate the diffuse emission in the F335M filter image from numerous point sources (mostly stars) and compact sources (mostly galaxies). We proceed in four main steps, which we outline here. More details are given in Appendix A.
We model and subtract the large diffraction spikes from two bright stars Merope and PQ Tau. These two stars are outside the JWST fields, but their spikes are prominent structures in Field 1. One of the Merope spikes is also visible in Field 2.
We subtract from a first approximation the diffuse emission using a multiscale median filter. This step yields background-subtracted images, which we then use to identify and remove point sources and compact sources located within the JWST fields.
We subtract point sources using an empirical template of the JWST point spread function (PSF). This template was obtained by stacking background-subtracted images centered on point sources to subpixel accuracy. Point sources are subtracted by fitting the PSF template over small images centered on each source. We do not use the core of the point sources in this fit because it is often saturated.
Compact extended sources — including galaxies and the saturated cores of bright stars — are identified on maps after background, spike, and point source subtraction, and are then masked. The sky area we mask is set by a brightness threshold applied to maps in which spikes, multi-scale background, and point sources have been subtracted (see Appendix A). The masked areas are filled in with the multi-scale background maps.
To assess systematic uncertainties associated with the data cleaning, we check how sensitive the results of our data analysis are to the brightness threshold used to create the mask. For the fiducial mask, 9% of the sky pixels are masked. We use two alternative masks, the light mask for which this fraction is 5% and the heavy mask for which it is 14%.
For statistical analysis, we divide the dither pointings into two sets and , which we use to produce a pair of images for each of the two fields. The two images in each pair cover the same sky region, but have independent statistical noise. Each of the four images (2 pairs) is processed separately by following steps (1) to (4). Separate background maps and masks are computed for each image. After step (4), the difference between the two images in each pair allows us to identify and mask a small number of glitches present in one of the two images. Mathematically, for each field, we obtain two independent maps of diffuse emission and from sky images and with spikes and stars subtracted:
@@
For power spectrum analysis, we cut out two square images, with 2000 pixels on each side, in each field. The positions of these cuts are drawn with solid black lines in the JWST images displayed in Fig. 1. The center positions and orientation of these images can be found in Table 1. We analyze eight square images (four pairs) from which diffraction spikes, stars, galaxies, and glitches have been removed. The maps of diffuse emission in each of the four pairs are presented in Fig. 2. The four images show a similar pattern with nearly parallel striations. The JWST/NIRCam pixel size corresponds to mpc (8.5 au) at the distance of the Pleiades. The full size of each square image is 80 mpc. Compared with other observational means, one may say that the JWST offers us a microscope view at the diffuse ISM.
@@
The mean brightness has been subtracted from the JWST images because it is a model value (Rigby et al., 2023) introduced in the data pipeline that does not separate interstellar emission from the zodiacal, stellar, and extragalactic backgrounds. To estimate the mean PAH emission, we measured the mean brightness averaged over each image field of view at m, using the Wide-Field Infrared Explorer (WISE) data (Wright et al., 2010). The values listed in Table 1 represent the emission from the Pleiades nebula with the large-scale Galactic contribution, measured off the nebula, subtracted out. The brightness ratios between the diffuse Galactic emission in the F335M NIRCam filter and the WISE m filter, values listed in Table 1, were measured by correlating the structure of the 12m emission with the multiscale background F335M maps. These values are consistent with dust emission models (Compiègne et al., 2011; Draine and Li, 2007) and observations222The color ratio is related to the size distribution of PAHs. The diffuse ISM values, derived from the analysis of the DIRBE and Spitzer/IRAC data, are in the range 2 to (Dwek et al., 1997; Flagey et al., 2006).
@@
Notes. (a) Coordinates of image centers. (b) Position angle of the long axis of the JWST mosaic with respect to the North direction measured positive towards East (counterclockwise direction). (c) Mean emission at m averaged over each of the Pleiades images, measured using the WISE full resolution image. The background contribution is subtracted. (d) Emission ratio derived from a linear correlation of the NIRCam F335M background image and the WISE data.
@@ 3.1 Computation and statistical uncertainties
We compute the two-dimensional cross-power spectrum for a pair of square images and introduced in Eq.LABEL:eq:im_painting.
where the wavenumber and the angle , measured from the Celestial North modulo , are polar coordinates in the Fourier plane. denotes the real part, the hat symbol the Fourier transform, and the asterisk the complex conjugate. represents the image apodization applied to avoid edge effects in the Fourier transform333We use the Tukey function in the SciPy python package with its default parameters..
We denote the mean value of , within bins of width and , and the cross-power spectrum averaged over angles. To estimate uncertainties on and , we use the following formulae :
where is the noise power spectrum of images and , is the number of Fourier modes in the bin , and is the effective area of the sky covered by the maps after apodization. We refer to Knox (1995) and Hivon et al. (2002) for a mathematical derivation and the numerical assessment of these formulae, and to Viero et al. (2013) and Córdova Rosado et al. (2024) for their application to the analysis of infrared sky maps. To apply this formula, we require the noise power spectra for each pair of square maps. These were estimated from the power spectrum of the difference , for Fields 1 and 2 separately. The factor 2 arising from the difference was taken into account.
@@
Figure 4 shows the cross power spectra of the four JWST images. Hereafter, the wavenumber expressed in units of in the bottom axis is converted into physical scales in au on the top axis for a distance to the Pleiades of 135 pc (Melis et al., 2014; Abramson, 2018). The spectra are corrected for the signal attenuation by the telescope beam, using the power spectrum of the point spread function (PSF) image produced with the STPSF Python package for our NIRCam observations in the F335M filter. The PSF varies across the field of the camera, but this effect is ignored because it is smeared out by the raster mapping of the sky. The spectra in Figure 4 all show a flattening at high . This is not due to data noise, because the noise is not correlated between maps and .
@@ 3.2 Anisotropy model
In order to analyze the anisotropy of the diffuse emission, we average into logarithmic bins of k with and linear bins of , modulo , with , which do not overlap. We denote and the bin centers and the binned value of . The uncertainty on is estimated using Eq. 3. We model the angular dependence of with a Gaussian profile. Mathematically our data model is as follows:
@@
The angle defines the orientation, perpendicular to the striations seen in the images, for which the power is maximum, and measures the dispersion of this orientation over the images. represents the ratio of maximum to minimum power while the minimum power within the bin. We perform a global fit of with values of and independent of . These two model parameters are listed in Table 2 for each of the four images in Fig. 2. The values of match the dust polarization angles measured at the image centers using Planck polarization data at 353 GHz (see Appendix B and Fig. 21). In emission, the dust polarization angle is perpendicular to the projection of the magnetic field on the plane of the sky. As measures the orientation perpendicular to the striations, we conclude that the filamentary structure of the PAH emission is aligned with the magnetic field projection. This is similar to what is observed on parsec scales with Planck data (Clark et al., 2015; Planck Collaboration, 2016c).
@@
Figure 5 presents the angular profiles of the power spectrum amplitude obtained for Cut 1 in Field 1. The profiles are shown for five -bins, equally spaced in log() from large to small angular scales from top to bottom; the dashed lines represent the model fit to the data. We note that the profile appears to be less anisotropic for high wavenumbers444The fit shown in Fig. 5 allows for small variations of and between the five bins..
@@
The model allows us to derive the power spectra for the orientations parallel and perpendicular to the striations in the sky image:
@@
where is the power spectrum of the JWST PSF. The division by corrects the spectra for the attenuation of the beam. The data error bars are propagated to and by the fitting procedure. Our definition of corresponds to the power spectrum of the emission parallel to the mean orientation of the striations555This is true provided that the Gaussian in Eq. 5 is negligible at of , a condition that is satisfied for the values of in Table 2.. As Cho and Vishniac (2000) pointed out in their analysis of MHD turbulence simulations, the Fourier transformation only partially measures the three-dimensional anisotropy of the PAH emission because the striations orientation varies across the sky and along the line of sight. The angular widths of the profiles in Fig. 5 reflect the dispersion of the orientation of the striations across the image. These widths are significant. They exceed the dispersion of the polarization angles () reported by Planck Collaboration (2016a) in their analysis of Planck data in the southern Galactic cap.
3.3 CIB residual

### v2

@@
Interstellar medium studies rely on magnetohydrodynamic turbulence as a framework for interpretation. In this context, the statistical characterization of interstellar observations is of prime importance. We present a new perspective on diffuse interstellar matter by analyzing James Webb Space Telescope (JWST) observations of the Pleiades nebula with NIRCam. These observations are remarkable in that they provide a ”microscope” view of the cold neutral medium (CNM) with a spatial resolution of 0.2 mpc (40 au). A 2D Fourier analysis was used to characterize the structure of polycyclic aromatic hydrocarbon (PAH) emission in regions near and far from the Pleiades star Merope. To produce maps of the interstellar emission, stars and galaxies were filtered out. The final step in the data cleaning involved subtracting a component, in Fourier space, which we inferred to be a residual of the near-infrared cosmic background. The PAH emission power spectra are highly anisotropic. PAH emission power spectra are highly anisotropic at all scales, with a constant level of power anisotropy. The magnetic field orientation, as derived from the Planck dust polarization data, aligns with the PAH anisotropy. The spectra are well fitted with a break-free power law, suggesting that we do not observe a specific scale for energy dissipation. Power-law indices are -3.5 near Merope and -3 in the more distant field. These findings are discussed in relation to interstellar turbulence that may be driven by the Pleiades stars. The JWST observations of the Pleiades offer a new viewpoint for comparing observations and theoretical models, as they examine physical scales at which turbulence in the CNM is subsonic and decoupled from the thermal instability. The observations may indicate that the turbulent energy cascade in the CNM is anisotropic.
@@
Studies of the cold interstellar medium (ISM) frequently rely on numerical simulations of magnetohydrodynamic (MHD) turbulence as a theoretical model (Hennebelle and Falgarone, 2012; Burkhart, 2021). In this context, the statistical analysis of interstellar observations is crucial to interpreting observations (Richard et al., 2025). This paper presents a new perspective, using observations of the Pleiades nebula by the James Webb Space Telescope (JWST) with the Near-Infrared Camera (NIRCam).
@@
The Pleiades reflection nebula is an iconic astronomical object familiar to a wide audience and also a prime scientific target for studies of the dusty ISM (Gibson and Nordsieck, 2003a, b; Gibson, 2007). We are observing a chance encounter between a star cluster and diffuse molecular gas (Gordon and Arny, 1984; Herbig and Simon, 2001; Ritchey et al., 2006). The brightest members of the Pleiades cluster are late-B stars at a distance of 135 pc (Abramson, 2018) illuminating their surrounding ISM over a few parsecs (Gibson and Nordsieck, 2003a). The stellar illumination enables the ISM structure to be observed at smaller angular scales than is usually possible. The reflection nebulosity offers a spectacular view of the filamentary structure of the diffuse ISM.
@@
Many studies have relied on dust continuum observations to determine the power spectrum of interstellar matter (Gautier et al., 1992; Ingalls et al., 2004; Miville-Deschênes et al., 2007, 2010, 2016). The advent of Planck made it possible to extend this type of analysis to polarized dust emission (Planck Collaboration, 2020a), which established the statistical alignment of diffuse interstellar structures with the magnetic field direction on parsec scales (Clark et al., 2015; Planck Collaboration, 2016c). This result has been interpreted as a hallmark of magnetized interstellar turbulence (Caldwell et al., 2017; Kandel et al., 2017). In a broader context, the JWST offers a unique opportunity. Its angular resolution and sensitivity to the emission of polycyclic aromatic hydrocarbons (PAHs) enable imaging of the cold ISM at finer scales than was previously possible. At the distance of the Pleiades, the angular resolution of NIRCam corresponds to a spatial scale of 0.2 mpc, at which we may be able to identify signatures of dissipation of turbulence, in particular through ion-neutral friction (Burkhart et al., 2015; Hu et al., 2024). Although there is observational evidence of structure in the cold neutral medium (CNM) at these scales, existing observations are too sparse to conduct a statistical study (Heiles, 1997; Stanimirović and Zweibel, 2018).
@@
The purpose of this work is to demonstrate and capitalize on the potential of JWST for diffuse ISM studies. We used PAH emission to investigate the structure of the cold magnetized ISM. In fact, PAHs are tightly coupled to the gas through collisions (Omont, 1986) and to magnetic field lines. In the physical conditions of the Pleiades nebula – specifically the gas density and temperature and the radiation field strength – PAHs exist partly as cations and partly in neutral form (Flagey et al., 2006). However, both charged and neutral PAHs are coupled to magnetic field lines because their charge state varies on timescales much shorter than their drag time through collisions with the gas (Lepp and Dalgarno, 1988).
@@
The paper is organized as follows. Section 2 introduces the JWST observations and outlines the data processing steps to generate images of diffuse emission. Our 2D Fourier analysis of these images is described in Sects. 3 and 4. In Sect. 5 we discuss the results of the data analysis in relation to interstellar turbulence. Finally, Sect. 6 summarizes our main results. The paper includes three appendices. Appendix A details and illustrates the methods used to remove stars and mask galaxies from the JWST images. Appendix B provides a broader context for the JWST observations, in which we characterize the magnetized interstellar matter associated with the Pleiades nebula using Planck data. In Appendix C we use the observations to characterize the physical conditions and demonstrate that the JWST observations of the Pleiades probe the structure of the CNM.
@@ 2.1 JWST observations
The NIRCam instrument (Rieke et al., 2023) was used to image two fields, each measuring by as part of the GO 2143 program. Field 1 is centered approximately 2′ to the east of the bright Pleiades star Merope (23 Tau), and Field 2 about 10′ to the southeast of Merope. The top images in Fig. 1 show the location of the two fields within the Pleiades. The two images at the bottom were produced by the JWST pipeline. For this paper, we used version 1.18.0 of the calibration software.
The two fields were imaged in the F212N filter, centered on the 1-0 S(1) H2 line, and the F335M filter that includes the 3.3-m emission bands from PAHs. We also obtained observations at both sky positions in the F200W and F277W filters. The observations in the short- and long wavelength channels are executed simultaneously, in the spectral filters F212N with F335M and F200W with F277W. For each pointing, the integration time was for the H2 and PAH observations, and for the wide filters. In this study, we only analyzed the F335M data. The data obtained using the three additional filters will be presented in a future paper.
The long channel of NIRCam comprises two square detectors, each with 2000 pixels square and a field of view of . We used the FULLBOX 6TIGHT111NIRCam page at https://jwst-docs.stsci.edu/jwst-near-infrared-camera/nircam-observing-modes/ dithering pattern to obtain multiple offset pointings of the camera and cover the rectangular field without gaps. The dithering also mitigates bad pixels and flat-field uncertainties; it significantly reduces the data noise on the smallest angular scales, where brightness variations are the smallest. We did not use secondary sub-pixel dithers because the fine sampling of the point-spread function (PSF) is not essential to our observations.
@@ 2.2 Maps of diffuse emission
In order to study the PAH emission, we needed to separate the diffuse emission in the F335M filter image from numerous point sources (mostly stars) and compact sources (mostly galaxies). We did this in four main steps, which we outline here. More details are given in Appendix A.
We modeled and subtracted the large diffraction spikes from two bright stars, Merope and PQ Tau. These two stars are outside the JWST fields, but their spikes are prominent structures in Field 1. One of the Merope spikes is also visible in Field 2.
We subtracted from a first approximation the diffuse emission using a multi-scale median filter. This step yields background-subtracted images, which we then used to identify and remove point sources and compact sources located within the JWST fields.
We subtracted point sources using an empirical template of the JWST PSF. This template was obtained by stacking background-subtracted images centered on point sources to sub-pixel accuracy. Point sources are subtracted by fitting the PSF template over small images centered on each source. We did not use the core of the point sources in this fit because it is often saturated.
Compact extended sources — including galaxies and the saturated cores of bright stars — are identified on maps after background, spike, and point source subtraction, and are then masked. The sky area we masked is set by a brightness threshold applied to maps in which spikes, multi-scale background, and point sources have been subtracted (see Appendix A). The masked areas are filled in with the multi-scale background maps.
To assess systematic uncertainties associated with the data cleaning, we checked how sensitive the results of our data analysis are to the brightness threshold used to create the mask. For the fiducial mask, 9% of the sky pixels are masked. We used two alternative masks, the light mask (for which this fraction is 5%) and the heavy mask (14%).
For statistical analysis, we divided the dither pointings into two sets and , which we used to produce a pair of images for each of the two fields. The two images in each pair cover the same sky region, but have independent statistical noise. Each of the four images (2 pairs) is processed separately by following steps (1) to (4). Separate background maps and masks are computed for each image. After step (4), the difference between the two images in each pair allows us to identify and mask a small number of glitches present in one of the two images. Mathematically, for each field, we obtained two independent maps of diffuse emission, and , from sky images and with spikes and stars subtracted:
@@
For power spectrum analysis, we cut out two square images, with 2000 pixels on each side, in each field. The positions of these cuts are drawn with solid black lines in the JWST images displayed in Fig. 1. The center positions and orientation of these images can be found in Table 1. We analyzed eight square images (four pairs) from which diffraction spikes, stars, galaxies, and glitches have been removed. The maps of diffuse emission in each of the four pairs are presented in Fig. 2. The four images show a similar pattern with nearly parallel striations. The JWST/NIRCam pixel size corresponds to mpc (8.5 au) at the distance of the Pleiades. The full size of each square image is 80 mpc. Compared with other observational means, one can say that the JWST offers us a ”microscope” view at the diffuse ISM.
@@
The mean brightness has been subtracted from the JWST images because it is a model value (Rigby et al., 2023) introduced in the data pipeline that does not separate interstellar emission from the zodiacal, stellar, and extragalactic backgrounds. To estimate the mean PAH emission, we measured the mean brightness averaged over each image field of view at m, using the Wide-Field Infrared Explorer (WISE) data (Wright et al., 2010). The values listed in Table 1 represent the emission from the Pleiades nebula with the large-scale Galactic contribution, measured off the nebula, subtracted out. The brightness ratios between the diffuse Galactic emission in the F335M NIRCam filter and the WISE m filter, values listed in Table 1, were measured by correlating the structure of the 12m emission with the multi-scale background F335M maps. These values are consistent with dust emission models (Compiègne et al., 2011; Draine and Li, 2007) and observations222The color ratio is related to the size distribution of PAHs. The diffuse ISM values, derived from the analysis of the DIRBE and Spitzer/IRAC data, are in the range 2 to (Dwek et al., 1997; Flagey et al., 2006).
@@ 3.1 Computation and statistical uncertainties
We computed the 2D cross-power spectrum, for the pair of square images and introduced in Eq. LABEL:eq:im_painting:
where the wavenumber and the angle , measured from the celestial north modulo , are polar coordinates in the Fourier plane. denotes the real part, the hat symbol the Fourier transform, and the asterisk the complex conjugate. represents the image apodization applied to avoid edge effects in the Fourier transform444We used the Tukey function in the SciPy python package with its default parameters..
We use to denote the mean value of , within bins of width and , and the cross-power spectrum averaged over angles. To estimate uncertainties on and , we used the following formulae:
where is the noise power spectrum of images and , is the number of Fourier modes in the bin , and is the effective area of the sky covered by the maps after apodization. We refer to Knox (1995) and Hivon et al. (2002) for a mathematical derivation and the numerical assessment of these formulae, and to Viero et al. (2013) and Córdova Rosado et al. (2024) for their application to the analysis of infrared sky maps. To apply this formula, we required the noise power spectra for each pair of square maps. These were estimated from the power spectrum of the difference , for Fields 1 and 2 separately. The factor 2 arising from the difference was taken into account.
@@
Figure 4 shows the cross-power spectra of the four JWST images. Hereafter, the wavenumber expressed in units of in the bottom axis is converted into physical scales in au on the top axis for a distance to the Pleiades of 135 pc (Melis et al., 2014; Abramson, 2018). The spectra are corrected for the signal attenuation by the telescope beam, using the power spectrum of the PSF image produced with the STPSF Python package for our NIRCam observations in the F335M filter. The PSF varies across the field of the camera, but this effect is ignored because it is smeared out by the raster mapping of the sky. The spectra in Fig. 4 all show a flattening at high . This is not due to data noise, because the noise is not correlated between maps and .
@@ 3.2 Anisotropy model
To analyze the anisotropy of the diffuse emission, we averaged into logarithmic bins of k with and linear bins of , modulo , with , which do not overlap. We use and to denote the bin centers and the binned value of . The uncertainty on is estimated using Eq. 3. We modeled the angular dependence of with a Gaussian profile. Mathematically our data model is as follows:
@@
The angle defines the orientation, perpendicular to the striations seen in the images, for which the power is maximum, and measures the dispersion of this orientation over the images. represents the ratio of maximum to minimum power while the minimum power within the bin. We performed a global fit of with values of and independent of . These two model parameters are listed in Table 2 for each of the four images in Fig. 2. The values of match the dust polarization angles measured at the image centers using Planck polarization data at 353 GHz (see Appendix B and Fig. 21). In emission, the dust polarization angle is perpendicular to the projection of the magnetic field on the plane of the sky. As measures the orientation perpendicular to the striations, we conclude that the filamentary structure of the PAH emission is aligned with the magnetic field projection. This is similar to what is observed on parsec scales with Planck data (Clark et al., 2015; Planck Collaboration, 2016c).
@@
Figure 5 presents the angular profiles of the power spectrum amplitude obtained for Cut 1 in Field 1. The profiles are shown for five -bins, equally spaced in log() from large to small angular scales from top to bottom; the dashed lines represent the model fit to the data. We note that the profile appears to be less anisotropic for high wavenumbers555The fit shown in Fig. 5 allows for small variations of and between the five bins..
@@
The model allowed us to derive the power spectra for the orientations parallel and perpendicular to the striations in the sky image:
@@
where is the power spectrum of the JWST PSF. The division by corrects the spectra for the attenuation of the beam. The data error bars are propagated to and by the fitting procedure. Our definition of corresponds to the power spectrum of the emission parallel to the mean orientation of the striations666This is true provided that the Gaussian in Eq. 5 is negligible at of , a condition that is satisfied for the values of in Table 2.. As Cho and Vishniac (2000) pointed out in their analysis of MHD turbulence simulations, the Fourier transformation only partially measures the 3D anisotropy of the PAH emission because the striations orientation varies across the sky and along the line of sight. The angular widths of the profiles in Fig. 5 reflect the dispersion of the orientation of the striations across the image. These widths are significant. They exceed the dispersion of the polarization angles () reported by Planck Collaboration (2016a) in their analysis of Planck data in the southern Galactic cap.
| Field | |||
| deg. | deg. | deg. | |
| (a) | (a) | (b) | |
| Field 1 | |||
| Cut 1 | |||
| Cut 2 | |||
| Field 2 | |||
| Cut 1 | |||
| Cut 2 |
3.3 Cosmic near-infrared background residual